40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 13 21:30:55 2020, shruti, Update, ALS, ALS OOL noise with arms locked 20200213_ALS.pdf
    Reply  Fri Feb 14 14:02:13 2020, shruti, Update, ALS, ALS OOL noise with arms locked 819472_10.pdf225642_10.pdf693500_10.pdf
       Reply  Fri Feb 14 14:52:41 2020, gautam, Update, ALS, ALS OOL noise with arms locked 
          Reply  Tue Feb 18 18:14:59 2020, shruti, Update, ALS, ALS OOL noise with arms locked ALSNoise.pdf
             Reply  Wed Feb 19 22:20:22 2020, rana, Update, ALS, ALS OOL noise with arms locked 
             Reply  Fri Feb 21 10:59:08 2020, shruti, Update, ALS, PDH error signals? IMG_0082.jpgIMG_0083.jpgIMG_0084.jpgIMG_0085.jpg
                Reply  Fri Feb 21 13:02:53 2020, Koji, Update, ALS, PDH error signals? 
                   Reply  Fri Feb 21 20:44:18 2020, shruti, Update, ALS, ALS OOL noise and PDH ALSNoise.pdfIMG_0086.jpg
                      Reply  Sun Feb 23 18:15:22 2020, rana, Update, ALS, ALS OOL noise and PDH 
Message ID: 15218     Entry time: Fri Feb 21 10:59:08 2020     In reply to: 15216     Reply to this: 15219
Author: shruti 
Type: Update 
Category: ALS 
Subject: PDH error signals? 
Here are a few PDH error signals measured without changing the servo gain or phase from that optimized for 231.25 kHz. This was done by keeping the X arm cavity and laser unlocked but keeping the shutter for green open; so I did not force a frequency sweep but saw the unhindered motion of cavity wrt the laser using the PDH servo error monitor channel from the box (not sure if this is the best way to do it?).
 
Koji mentioned that there is a low pass filter with a cutoff frequency probably lower than 700 kHz which at the moment would hinder the efficacy of the locking at higher frequencies. The transfer function on the wiki suggests the same, although we are yet to investigate the circuit.
 
I measured the maximum range in the linear region of the signal, and here are the results:
  • Attachment 1: 231.25 kHz (current PDH sideband mod freq): 1.7 V
  • Attachment 2: 225.642 kHz: 1.2 V
  • Attachment 3: 100 kHz: 900 mV
  • Attachment 4: 763.673 kHz: 220 mV
Right now we have only inverted the phase to try locking at different frequencies (no finer adjustments were performed so elog 15216 cannot be an accurate representation of the true performance)
 
Ideas from the 40m meeting for adjusting the phase:
  1. Delay line for adding extra phase (would require over 40m of cable for 90 deg phase shift)
  2. Using two function generators for generating the sideband, clocked to each other, so that one can be sent to the PZT and the other to the mixer for demodulation.
  3. Use a different LPF (does not seem very useful for investigating multiple possible frequencies)

Once we adjust the phase we may be able to increase the servo gain for optimal locking. Unless it may be a good idea to increase the gain without optimizing the phase?

Attachment 1: IMG_0082.jpg  482 kB  | Hide | Hide all
IMG_0082.jpg
Attachment 2: IMG_0083.jpg  499 kB  | Hide | Hide all
IMG_0083.jpg
Attachment 3: IMG_0084.jpg  545 kB  | Hide | Hide all
IMG_0084.jpg
Attachment 4: IMG_0085.jpg  539 kB  | Hide | Hide all
IMG_0085.jpg
ELOG V3.1.3-